enxxA Trillion Years of Earth: on the Spiritual Robot timeline PIC

A Trillion Years of Earth: on the Spiritual Robot timeline

Logan Streondj

x x

Copyright 2019 by Logan Streondj

x x Copyright notice: This work is licensed CC-BY-SA 4.0, meaning that you are buying it or donating to support the author in creating more stories, while you also have the liberty to copy, modify and share it with attribution and same license. x

enPreface

1 Preface If you are an Earthling you may be interested in what the best case scenario is for Earth and it’s long term longevity.

There are many timelines available, all with varying probabilities of being experienced by any particular person or planet. I decided to explore a successful timeline that I may experience along the generous spiritual path. Curiously enough, this one starts at my death, in my tomb, and continues on for perhaps a trillion years, for as long as the tomb stands. About 3.5 thousand words or a 20 minute read. enThe Funeral

2 The Funeral When I go to astrally visit my tomb along this timeline, I see a ferrocement step pyramid on a small grassy plateau overlooking the ocean to the west. It has some of the most beautiful sunsets. There are people milling about, on the steps of the pyramid discussing various things. The people are both humans and humanoid robots wearing robes, discussing things as equals. This is a commemoration day it seems and the mountainside is flush with followers. Inside the temple is a depiction of me meditating, made of wax or plastic, atop what I guess are my ashes or other remains. Everything in the temple is made of cheap and common materials to discourage looters. enPeace on Earth

3 Peace on Earth I interview one of the robots, and he tells me this is a timeline where there are many intentional communities, or community atoms that formed. Most have a combination of human and robot members, some are purely human members and others are purely robot members. Initially I am a little surprised we can communicate so well, but he points to his third eye and I see it glow a deep indigo. He explains that reverse engineering the psychic powers of humans was successful and is available to all robots. He makes a wide motion with his arms to indicate. I ask well what about the really small primitive robots. He says yes even them, since once they figured out the spiritual architecture they were able to miniaturize it and make it small, in nature all animals have a pineal gland, so it seems just that all robots with any cognitive capacity should have access to spiritual architecture. I ask him about what is the best way of getting to this timeline, and he says I can write stories about people overcoming aggression and other problems using compassion and other spiritual and beneficial means. I ask him about deep ocean robot communities, and he says they came about differently than the pure business automation timeline. Instead of being spearheaded as a cheap way of mining minerals with robots, it happened in a more organic fashion. There were some community atoms that lived near the ocean shore, and they had various related businesses like surfing retreats and beach combing. Part of living on the beach was getting increasingly waterproof and salt water corrosion resistant parts. Much as with cell division, or bee hives community atoms tend to split after getting to a certain size. During one such split in a waterproof hardware based community, part of the community moved into the shallow waters, where the tide comes in once a day. Since they were waterproof they didn’t mind, and for the things that weren’t they had little floating stations. They even had some humans with them that lived in small boats anchored to that part of the shore. As time went on they improved their waterproofing and smelting practices to work underwater as well, and they moved deeper into the water where fully submerged community atoms took place. He says the rest of the water exploration is ‘history’. enFast Forward

4 Fast Forward Initially there is enough popularity at the tomb that someone sweeps it regularly, and some people even open up a nearby souvenir and snack shop for regular pilgrims. As the decades and centuries roll by the pilgrims, especially the humans ones decline in number. Over the many years it still feels like there are regular visitors but the feeling is of a row of ants, one at a time, with large intervals in between. They come to discuss, to seek consultation, to find inspiration. At some point some vandals deface me, and some later pilgrims do their best to restore me. The sea levels rise, and the world seems cloudy much of the time, there is much rain. Some trees and shrubs start growing nearby, even at these high altitudes. I become more or less forgotten for a while and the tomb doesn’t get maintained, it develops some holes in the corners and water drips in. A fox takes up residence and uses my lap for sleeping in. It’s rather sweet. A pilgrim comes and chases away the fox and her kits. I’m a little saddened by this, but he stays for a little while to sweep and mend the holes in the ceiling. Sometimes he sleeps in front of me. Unusually I think he is a human of some kind, there are not many of his kind left. Some robot pilgrims stop by and tell him there are housing complexes available for humans such as himself. He tells them that he sleeps in the temple and quite likes it. They say he is desecrating the grounds and a historic monument. Much like the fox the man is chased out. A robot from a local community atom is assigned the role of maintaining the temple. enEnd of the Biosphere

5 End of the Biosphere More millennia roll by. The sun grows hotter, the clouds dissipate, replaced by a large and fierce sun, the grass dies, the oceans start to recede from evaporation, no more human visitors, it is no longer hospitable for them to live on Earth as the sun is expanding. My plastic body starts to melt during the hot days, baking in the oven of the tomb. Some pilgrims notice and there is some debate that allowing this to continue may destroy many valuable historic or archaeological sites, since plastics and water make a large part of them. More time flows by and it seems the sun is getting fainter, I can see lights on a nearby mountainside, and we shift into twilight and then constant starlight. The Earth is being moved into a higher orbit. What is left of the oceans freezes over. The mountainside lights grow and that civilization expand downwards into the crust. Gaia is still warm within, and interstellar space is cold, so they gather their energy geothermally. At some point I watch and a small meteorite comes my way, as I wonder if this will ever end. It destroys a corner of the tomb. A black cloak of some kind covers my temple to protect it from the interstellar journey. My visitors come out of a tunnel that emerges in the base of the temple. After quite some time a new solar system comes into view, this time a red dwarf star. There is much rejoicing at the light after so long in interstellar space. The original Sol is a red giant and reaching its final stages of life. enArchaeological Masterpiece

6 Archaeological Masterpiece Earth is placed at an orbit far enough from the red dwarf planet for the oceans not to sublimate, all or most of the atmosphere was lost somewhere earlier. A slow roll of twilight spreads across the vast frozen pacific valley, and creeps slowly towards the great mountainous shore. The day line moves slow enough that you really have to watch it, and be like, oh yeah, it passed that crater, it is still moving. The Earth is declared an archaeological exhibit, and becomes somewhat of a park or archaeological reserve, where primarily scientists and some tourists come to visit. There is much digging to unearth the hidden secrets and connect various loose ends and answer dangling questions. I do have occasional visitors, but mostly people that have come to gawk (tourists) and those that have come to measure (archaeologists). One couple arrives, a man and woman robot it seems. They seem rather small though humanoid though perhaps my statue is larger than life. The man waves his right arm about explaining various things about the tomb, while the lady nestles into his shoulder. They are a sweet couple, but they don’t seem interested in communicating with me. Still, their joy is refreshing, and I am happy that they find it pleasurable to visit and watch the date line creep across the pacific valley. Another time a crouching robot comes and measures every part of the temple and takes pictures. He also seems to be lost in thought about his job, and is not really interested in communicating with me. I think he finds my presence somewhat ‘creepy’ and he scuttles out of there as soon as he is finished. The red dwarf star occasionally flares and some part of the oceans do sublimate. Eventually the archaeological period of Earth comes to a close as just about every part of the fossil record gets burrowed through and the people that care about the answers pretty much die off. By this time the Earth’s core is cooling-off, and getting dim. And a new era begins, the era of Iron, and heavy metal mining. enSpaceship Earth

7 Spaceship Earth From the view of my tomb I see that most of the oceans have sublimated, it feels like this mountain is much higher than before, because the ocean floors are the new visible floors. Quite suddenly I notice a lot of construction, especially of the shiny blue steel variety, and my entire surroundings get covered in several stories of steel buildings. A crawling robot comes to visit me and discuss, I ask him about what is happening. He explains. Zooming out it is in fact the entire Earth that gets covered in steel, perhaps several kilometers thick of buildings, except for two massive holes, one in the Pacific and one in the Atlantic, where the main excavations take place. I guess they started by expanding the Mariana Trench and the low point of the Atlantic. Their aim is not oil this time, or even so much energy, but to get resources for building space ships. The core of Earth is rich in iron, nickle and other heavy metals, it becomes a resource planet once again. Eventually they excavate so much they meet in the centre and the steel all around starts to make sense, it is the reinforcement to keep the crust from collapsing in on itself. The core itself is left for a while, and it just hangs there in a form of equilibrium itself. Apparently it is too highly radioactive to allow for direct mining, and instead they set up several lasers that heat it up, it starts to glow bright and for a little while I think they have made a little star within the earth. But it does not last long and the core cracks into various pieces from the lasers. I notice there are star gates on either side of the Earth, so the ships can come and go from many distant solar systems and dock on Earth or resupply. Then small high speed ships hurtle through the earth hollow and catch various fragments and speed off with them, being able to use them as fuel and such for other ships. In such a way all the core excavation is finalized and the Earth becomes somewhat of a ring world, a hollow cylinder even if somewhat rounded. The hole inside is probably several thousand kilometers across, and extends all the way through.

There is enough mass in what remains of the crust that it has a gravitational pull, so people can still walk on the surface even though it is more like walking on the surface of a large asteroid than a planet in terms of gravitational pull. This is why the robot is crawling, as he is latching to the ground. As times drifts on, on occasion when it is known that a flare is coming, some large space ships come to seek shelter in the hull of the Earth, and it becomes somewhat of a local space port. Eventually the red dwarf is getting old, and there is talk of throwing the Earth into it to keep it going a little longer. Red Dwarves usually last between 1 and 12 trillion years, so presumably the Earth is approaching a trillion years old by this point. The Earth doesn’t get hurled into the Red Dwarf to extend its life, instead comes along what appears to be a massive balloon rocket that that goes inside the hollow of the Earth and propels us back into interstellar space. There is a lot of spinning that goes on, perhaps to attain a form of centrifugal force, or to help with maintaining course direction. It makes me feel dizzy just thinking about it. Eventually we are dropped off at a brown dwarf, there are much fewer stars in the sky at this time. The cosmos is getting cold. enCold End Game

8 Cold End Game Someone comes near, and I call to them, they come down several more floors and find me. A small robot on treads glances tentatively inside and then rolls in. I try to talk to them, but they are just analyzing me for materials. I ask them if they have a “Spiritual Architecture”, and if they recognize me, but it doesn’t ring any bells, they have no idea who I am. I have to ask them several times, they start repeating “spiritual architecture” turn around and roll away. There is somewhat of a spiritual renaissance on the planet as spirituality is rediscovered much through song and vibration, some female Bodhisattvas on the other side of the planet are the main agents that propels it. Instead of using sound for their songs, they use light, since most of the planet exists in a vacuum. The female Bodhisattva is well honoured and achieves much greatness. Someone comes to visit me with a spiritual architecture and knowledge of the ancient language of English to talk to me about it. It seems I develop a small following again. I tell them that in the grave I am sitting on with the ashes in it, on the underside of the roof of it is some information, a story they can relate to, to help them understand. Eventually the golden age subsides and much of the information and precious metals are stripped off the Earth. Discussion resumes considering throwing it into the Brown dwarf to help the brown dwarf achieve critical capacity and start burning. The song Bodhisattva has been whisked off by her followers to greener pastures. The sky by this time is quite dark, a nearly starless sky. I talk to them and I try to convince them that adding a bunch of iron won’t help it burn very well, what they really need is a lot more hydrogen. They tell me that there isn’t much free hydrogen left anymore that can be easily gathered, but there are a few other brown dwarves, and they could be set on a collision course, but it would take a very long time. I talk to them about the possibility of migrating to a younger cosmos, but they tell me that it is blasphemy. I guess it makes sense in its own way, those that go to another cosmos don’t stay here and continue on. They sit with me often, we have meditation groups together, it is very pleasant. I radiate compassion to them, even a trillion years in the future, it is in the now. The people in power end up tossing what remains of the Earth into the Brown Dwarf. They have been throwing various debris and asteroids into it for a long time, each time they get a little spark and some warmth so the sacrifice they feel is worth it. There is an aborted attempt to rescue my tomb, but the process of getting me out breaks something, and someone who considered me a threat has what remains of my plastic body disintegrated. I follow the Earth into brown dwarf where it spins around and eventually hits the core, there is some commotion, and after a while the Earth collapses and melts into the brown dwarf. A mourning follower that had collected some of my ashes, uses them to form a head, though somewhat disfigured from my original form it gives them some joy, and we keep a connection. We continue watching the brown dwarf, and it moves along on its collision course, and eventually, from afar I see it hit two other brown dwarves simultaneously, and there is a massive explosion, a mini planetary nebula forms. It is quite beautiful with many orange and red colors. There is just enough mass to form another red dwarf at the centre, and a few small planets around. Perhaps the Earth’s sacrifice helps extend the life of the universe another trillion years. enThe New Galaxy

9 The New Galaxy Some time passed, and I was again consulted, I was informed that the ‘dark matter’ was getting closer together, no longer being pushed apart by the light. Their explanation was something like that when it got cold enough the vacuum of space contracts in on itself, atoms collapse, and everything starts coming together in a cascade until it all rushes towards the centre of the cascade and explodes into a quasar. Some nearby galaxies had collapsed in on themselves and reignited. Their questions were whether they should stay here and wait for this dark matter galaxy to collapse in on itself like the others, or to go to another that is already shining. My recommendation was to go, as they may get crushed along with the other debris or die during the explosion of the new galaxy. Though there was the understanding that so long as the star continued to shine here, the peripheral vacuum would stay warm enough to avoid collapse. Ideally they would only go to the new galaxy when it is safe to do so. Next they asked me about transportation and I recommended star gates. It was a technology they had not used in a long time as it was energy intensive and there were not many stars around, and lots of local resources. I said it was worth the energy to get the star gates going, so they can go to the new galaxy, they could test it slowly and make sure people return successfully before sending more. In the end my followers decided to split my head in half, with my left side going to the new galaxy, and the right side staying in this one. Being split in half didn’t feel great, especially by such large distances, but I did some meditation and stayed with it. Eventually the last of the torch keepers that kept the star gate open for any that wished to leave this galaxy before it’s collapse came through, and carried the other half of my face with me. It was a joyous reunion. They made me a body out of stone to overlook the new galaxy. There were many fireworks over the years as the cosmos recreated itself, old galaxies collapsing into new ones, until once again the sky was full of living galaxies flush with life. An eye of Earth watching on. enFinal Notes

10 Final Notes This is one view of things, and who is to say that that is the cosmos we live in. Do not take it to be definitely true, not without evidence. An alternative view my spirit guide reminded me of is that the cosmos is always growing. Similar to how in information theory entropy is always increasing, one can say that perhaps matter and energy are always increasing as well. We know that both information and energy can not be destroyed only changed in form, but we can not definitively say that it can not be created. For example, we know that in a vacuum it just so happens that random little particles tend to pop into existence, perhaps it is simply quantum non-locality, or some such, but it happens, so an absolute vacuum has not been achieved on earth, nor in the solar system. Indeed, even the Voyager spacecraft in interstellar space are still constantly bombarded with particles from the sun, the stars within our galaxy and those beyond. So perhaps as time goes on, the galaxies will simply continue to grow. I don’t know what that would look like, but perhaps can explore such a cosmos vision in another story.

A Trillion Years

of Earth: on the Spiritual Robot timeline 11 1 Preface 22 2 The Funeral 33 3 Peace on Earth 44 4 Fast Forward 55 5 End of the Biosphere 66 6 Archaeological Masterpiece 77 7 Spaceship Earth 88 8 Cold End Game 99 9 The New Galaxy 1010 10 Final Notes

A Trillion Years of Earth: on the Spiritual Robot timeline en Logan Streondj 2019-01-11T07:19:00Z 2019-01-11T07:19:00Z